# Finite Sample Identification of Partially Observed Bilinear Dynamical Systems

### v1

@@ 1 Introduction
Notations: We use boldface lowercase (uppercase) letters to denote vectors (matrices). $\rho({\bm{X}})$, $\|{\bm{X}}\|_{\textup{op}}$ and $\|{{\bm{X}}}\|_{F}$ denote the spectral radius, spectral norm and Frobenius norm of a matrix ${\bm{X}}$, respectively. $\|{{\bm{v}}}\|_{\ell_{2}}$ denotes the Euclidean norm of a vector ${\bm{v}}$, and $({\bm{v}})_{i}$ denotes its $i$-th element. For a positive definite matrix ${\bm{M}}\in\mathbb{R}^{d\times d}$, the Mahalanobis norm of a vector ${\bm{v}}\in\mathbb{R}^{d}$ is given by $\|{{\bm{v}}}\|_{{\bm{M}}}=\sqrt{{\bm{v}}^{\top}{\bm{M}}{\bm{v}}}$. For a sequence of $d\times d$ matrices ${\bm{M}}_{1},\dots,{\bm{M}}_{k}$, we use the convention that $\prod_{i=1}^{k}{\bm{M}}_{i}={\bm{M}}_{1}\times{\bm{M}}_{2}\times\cdots\times{\bm{M}}_{k}$. We denote by $\mathcal{S}^{p-1}$, the centered unit sphere in $\mathbb{R}^{p}$. We denote by $a\vee b$, the maximum of two scalars $a$ and $b$. We use $\gtrsim$ and $\lesssim$ for inequalities that hold up to an absolute constant factor. $\tilde{\mathcal{O}}(\cdot)$ and $\tilde{\Omega}(\cdot)$ are used to show the dependence on a specific quantity of interest (up to constants and logarithmic factors). Finally, $\otimes$ denotes the Kronecker product.
@@ 2.2 Input Choice & Stability of Bilinear Dynamical Systems
Let $\mathcal{U}\subseteq\mathbb{R}^{p}$, $\kappa\geq 1$, and $10<\rho<$ 1. We say that a partially observed bilinear dynamical system (as defined in (2.1)) with state-transition matrices ${\bm{A}}:=\{{\bm{A}}_{0},\dots,{\bm{A}}_{p}\}$ is $(\mathcal{U},\kappa,\rho)$-uniformly-stable, if the joint spectral radius of the set $\mathcal{U}\circ{\bm{A}}$ satisfies: (i) $\rho(\mathcal{U}\circ{\bm{A}})<\rho<1$; and (ii) $\phi(\mathcal{U}\circ{\bm{A}},\rho)\leq\kappa$.
Observe that if there exists a nonempty and bounded set $\mathcal{U}\subseteq\mathbb{R}^{p}$ such that $\rho(\mathcal{U}\circ{\bm{A}})<1$, then for any $\rho(\mathcal{U}\circ{\bm{A}})<\rho<1$, the system is $(\mathcal{U},\kappa,\rho)$-uniformly-stable with $\kappa=\phi(\mathcal{U}\circ{\bm{A}},\rho)\vee 1$. We provide detailed discussion on this claim in Appendix A. Furthermore, we note that Definition 1 naturally generalizes that introduced by [35]. Indeed, there the authors assume that there exists ${\bm{u}}^{\star}$ such that $\rho({\bm{u}}^{\star}\circ{\bm{A}})<1$. This is equivalent to assuming that their system is ($(\{{\bm{u}}^{\star}\},\kappa,\rho)$-uniformly-stable for some $\kappa\geq 1$ and $\rho({\bm{u}}^{\star}\circ{\bm{A}})<\rho<1$. As we shall make precise shortly, we need stronger requirements on the stability of the system in comparison with [35] because we are concerned with the task of identification. This requirement stems from the need to have persistence of excitation so that estimation is possible.
@@ 3 Learning the Markov-like Parameters
$\bar{{\bm{u}}}_{t}^{\top}$ $:=\begin{bmatrix}1&{\bm{u}}_{t}^{\top}\end{bmatrix},\qquad\tilde{{\bm{u}}}_{t}:=\begin{bmatrix}{\bm{u}}_{t}\\ {\bm{u}}_{t-1}\\ \bar{{\bm{u}}}_{t-1}\otimes{\bm{u}}_{t-2}\\ \bar{{\bm{u}}}_{t-1}\otimes\bar{{\bm{u}}}_{t-2}\otimes{\bm{u}}_{t-3}\\ \vdots\\ \bar{{\bm{u}}}_{t-1}\otimes\bar{{\bm{u}}}_{t-2}\otimes\cdots\otimes{\bm{u}}_{t-L}\end{bmatrix},\qquad\tilde{{\bm{w}}}_{t}:=\begin{bmatrix}{\bm{z}}_{t}\\ {\bm{w}}_{t-1}\\ {\bm{w}}_{t-2}\\ {\bm{w}}_{t-3}\\ \vdots\\ {\bm{w}}_{t-L}\end{bmatrix},$ (3.2)
$\text{and}\quad{\bm{F}}$ $:=\begin{bmatrix}{\bm{I}}_{m}&{\bm{C}}&{\bm{C}}({\bm{u}}_{t-1}\circ{\bm{A}})&{\bm{C}}\prod_{\ell=1}^{2}({\bm{u}}_{t-\ell}\circ{\bm{A}})&\ldots&{\bm{C}}\prod_{\ell=1}^{L-1}({\bm{u}}_{t-\ell}\circ{\bm{A}})\end{bmatrix}.$
@@
${\bm{G}}$ $:=\left[\begin{array}[]{c c c c c}{\bm{D}}&{\bm{G}}_{1}&{\bm{G}}_{2}&\cdots&{\bm{G}}_{L}\end{array}\right]\in\mathbb{R}^{m\times d_{\tilde{u}}},\quad\text{with}\quad d_{\tilde{u}}=(p+1)^{L}+p-1,$ (3.3)
@@
With the dynamics written in the form of (3.4), it is natural to use the least squares estimation method to learn the Markov-like parameters from the observations $\{{\bm{y}}_{t},{\bm{u}}_{t}\}_{t=1}^{T}$ More specifically, the (minimum norm) Least Squares Estimator (LSE) $\hat{{\bm{G}}}$ of ${\bm{G}}$ admits a closed form and can be defined as:
@@
[…]
@@ 5 Numerical Experiments
For our experiments, we consider a partially observed bilinear dynamical system (2.1) with $n{=}5$ hidden states, input dimension $p{=}2$, and output dimension $m{=}2$. The dynamics matrices ${\bm{A}}_{0},{\bm{A}}_{1},{\bm{A}}_{2}$ are constructed with i.i.d. $\mathcal{N}(0,1)$ entries, and are scaled to have spectral radius $\rho({\bm{A}}_{0}){=}\rho_{0}$ and $\rho({\bm{A}}_{1}){=}\rho({\bm{A}}_{2}){=}\rho_{k}$, where $\rho_{0},\rho_{k}$ are hyper-parameters in our experiments. Similarly, ${\bm{B}},{\bm{C}}$ and ${\bm{D}}$ are generated with i.i.d. $\mathcal{N}(0,1/n)$ and $\mathcal{N}(0,1/m)$ entries, respectively. The noise processes $\{{\bm{w}}_{t}\}_{t=0}^{T}$, and $\{{\bm{z}}_{t}\}_{t=0}^{T}$ are chosen according to i.i.d. Gaussian distribution with zero mean and variances $\sigma^{2}=0.0001$. Lastly, the control inputs are either sampled uniformly at random from the sphere $\sqrt{p}\cdot\mathcal{S}^{p-1}$ or sampled i.i.d. from a Gaussian distribution $\mathcal{N}(0,{\bm{I}}_{p})$.
[…]
We observe double descent curves [38] for $L{=}7$. This is because our regression problem is unregularized and has $(p+1)^{L}+p-1=2188$ unknown parameters, and the number of input-features ($\tilde{{\bm{u}}}_{t}$) is $T-7$. Hence, for $L=7$, we see the peak at $T=2195$, and the error decays smoothly after this point. Note that the peak occurs at $T-L=(p+1)^{L}+p-1$ (where the number of unknown parameters become equal to the number of input-features). For $L=\{5,6\}$, we do not see the double descent because we start at $T=1000$ which is greater than $(p+1)^{L}+p-1$.
[…]
@@ B.2 Hypercontractivity to Bounded Moments (Covariates)
Isotropic Covariates: To begin, note that $\operatorname{\mathbb{E}}[\tilde{{\bm{u}}}_{t}]=0$ due to Assumption 2. Next, we will show that ${\bm{\Sigma}}[\tilde{{\bm{u}}}_{t}]=\operatorname{\mathbb{E}}[\tilde{{\bm{u}}}_{t}\tilde{{\bm{u}}}_{t}^{\top}]={\bm{I}}_{d_{\tilde{u}}}$ as follows: Let $[\tilde{{\bm{u}}}_{t}]_{0}={\bm{u}}_{t}$, and let $[\tilde{{\bm{u}}}_{t}]_{i}$ be the $i$-th partition of $\tilde{{\bm{u}}}_{t}$ with $p(p+1)^{i-1}$ entries for $i=1,\dots,L$. Then, we have that $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{0}[\tilde{{\bm{u}}}_{t}]_{0}^{\top}\right]=\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{1}[\tilde{{\bm{u}}}_{t}]_{1}^{\top}\right]={\bm{I}}_{p}$. Similarly, for $i=2,\dots,L$, we have
@@
for all $2\leq i<j$. Similarly, we also have $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{1}[\tilde{{\bm{u}}}_{t}]_{j}^{\top}\right]=\operatorname{\mathbb{E}}\left[{\bm{u}}_{t-1}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes{\bm{u}}_{t-j})^{\top}\right]=0$, for all $j=2,\dots,L$. Lastly, $[\tilde{{\bm{u}}}_{t}]_{0}$ is independent of $[\tilde{{\bm{u}}}_{t}]_{j}$, for all $j=1,\dots,L$, due to Assumption 2, hence, $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{0}[\tilde{{\bm{u}}}_{t}]_{j}^{\top}\right]=0$. Combining (B.2) and (B.3), we get
@@
$\bullet$ Quadratic Terms II: For $i\neq j$, consider the expectation $\operatorname{\mathbb{E}}\left[({\bm{v}}_{i}^{\top}[\tilde{{\bm{u}}}_{t}]_{i})^{2}({\bm{v}}_{j}^{\top}[\tilde{{\bm{u}}}_{t}]_{j})^{2}\right]$. In order to upper bound this term, we first upper bound an intermediate term as follows: Let $[\tilde{\tilde{{\bm{u}}}}_{t}]_{k}$ denote the $k$-th partition of $\tilde{\tilde{{\bm{u}}}}_{t}$ defined in (B.1), for $k=1,\dots,L$. Let ${\bm{q}}_{k},{\bm{q}}^{\prime}_{k}$ be two arbitrary vectors in ${\bm{R}}^{(p+1)^{k}}$. We will use the mathematical induction to prove the following intermediate result under Assumption 2,
@@
=$\qquad\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\left(q_{11}+{\bm{q}}_{12}^{\top}{\bm{u}}_{t-1}\right)^{2}\left(q_{11}^{\prime}+{\bm{q}}_{12}^{\prime\top}{\bm{u}}_{t-1}\right)^{2}\right],$
@@
=$\qquad\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\big{[}\operatorname{\mathbb{E}}\big{[}\left({\bm{u}}_{t-i}^{\top}{\bm{V}}_{i}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes\bar{{\bm{u}}}_{t-i+1})\right)^{2}\left((\bar{{\bm{u}}}_{t-i}^{\top}\otimes\cdots\otimes{\bm{u}}_{t-j}^{\top}){\bm{V}}_{j}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes\bar{{\bm{u}}}_{t-i+1})\right)^{2}$
@@
=$\qquad\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{k=1}^{(p+1)^{i-1}}\operatorname{\mathbb{E}}\big{[}\|{{\bm{V}}_{i}^{\top}{\bm{u}}_{t-i}}\|_{\ell_{2}}^{2}\big{(}{\bm{v}}_{j2k}^{\top}({\bm{u}}_{t-i}\otimes\bar{{\bm{u}}}_{t-i-1}\otimes\cdots\otimes{\bm{u}}_{t-j})\big{)}^{2}\big{]},$
=$\qquad\stackrel{{\scriptstyle\text{(b)}}}{{=}}\sum_{k=1}^{(p+1)^{i-1}}\operatorname{\mathbb{E}}\big{[}\|{{\bm{V}}_{i}^{\top}{\bm{u}}_{t-i}}\|_{\ell_{2}}^{2}\big{(}(\bar{{\bm{u}}}_{t-i-1}\otimes\cdots\otimes{\bm{u}}_{t-j})^{\top}{\bm{V}}_{j2k}{\bm{u}}_{t-i}\big{)}^{2}\big{]},$
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[({\bm{v}}_{1}^{\top}{\bm{u}}_{t-1})^{2}\|{{\bm{v}}_{j1}+{\bm{V}}_{j2}{\bm{u}}_{t-1}}\|_{\ell_{2}}^{2}\right],$
@@
$\bullet$ Quartic Terms: Consider the expectation $\operatorname{\mathbb{E}}\left[({\bm{v}}_{i}^{\top}[\tilde{{\bm{u}}}_{t}]_{i})^{4}\right]$. In order to upper bound this term, we first upper bound an intermediate term as follows: Recall the definition of $\tilde{\tilde{{\bm{u}}}}_{t}$ from (B.1), and let $[\tilde{\tilde{{\bm{u}}}}_{t}]_{k}$ denote its $k$-th partition with $(p+1)^{k}$ entries, for $k=1,\dots,L$. For any ${\bm{q}}_{k}\in{\bm{R}}^{(p+1)^{k}}$, we will use the mathematical induction to prove that
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\left(\bar{{\bm{u}}}_{t-k}^{\top}{\bm{Q}}_{k}[\tilde{\tilde{{\bm{u}}}}_{t}]_{k-1}\right)^{4}{~{}\big{|}~{}}{\bm{u}}_{t-k}\right]\right],$
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\left({\bm{u}}_{t-k}^{\top}{\bm{V}}_{k}[\tilde{\tilde{{\bm{u}}}}_{t}]_{k-1}\right)^{4}{~{}\big{|}~{}}{\bm{u}}_{t-k}\right]\right],$

### v2

@@ 1 Introduction
[…]
@@ 2.2 Input Choice & Stability of Bilinear Dynamical Systems
Let $\mathcal{U}\subseteq\mathbb{R}^{p}$, $\kappa\geq 1$, and $0<\rho<1$. We say that a partially observed bilinear dynamical system (as defined in (2.1)) with state-transition matrices ${\bm{A}}:=\{{\bm{A}}_{0},\dots,{\bm{A}}_{p}\}$ is $(\mathcal{U},\kappa,\rho)$-uniformly-stable, if the joint spectral radius of the set $\mathcal{U}\circ{\bm{A}}$ satisfies: (i) $\rho(\mathcal{U}\circ{\bm{A}})<\rho<1$; and (ii) $\phi(\mathcal{U}\circ{\bm{A}},\rho)\leq\kappa$.
Observe that if there exists a nonempty and bounded set $\mathcal{U}\subseteq\mathbb{R}^{p}$ such that $\rho(\mathcal{U}\circ{\bm{A}})<1$, then for any $\rho(\mathcal{U}\circ{\bm{A}})<\rho<1$, the system is $(\mathcal{U},\kappa,\rho)$-uniformly-stable with $\kappa=\phi(\mathcal{U}\circ{\bm{A}},\rho)\vee 1$. We provide detailed discussion on this claim in Appendix A. Furthermore, we note that Definition 1 naturally generalizes that introduced by [35]. Indeed, there the authors assume that there exists $\bm{u}$$^{\star}$ such that $\rho({\bm{u}}^{\star}\circ{\bm{A}})<1$. This is equivalent to assuming that their system is $(\{{\bm{u}}^{\star}\},\kappa,\rho)$-uniformly-stable for some $\kappa\geq 1$ and $\rho({\bm{u}}^{\star}\circ{\bm{A}})<\rho<1$. As we shall make precise shortly, we need stronger requirements on the stability of the system in comparison with [35] because we are concerned with the task of identification. This requirement stems from the need to have persistence of excitation so that estimation is possible.
@@ 3 Learning the Markov-like Parameters
$\bar{{\bm{u}}}_{t}^{\top}$ $=\begin{bmatrix}1&{\bm{u}}_{t}^{\top}\end{bmatrix},\qquad\tilde{{\bm{u}}}_{t}=\begin{bmatrix}{\bm{u}}_{t}\\ {\bm{u}}_{t-1}\\ \bar{{\bm{u}}}_{t-1}\otimes{\bm{u}}_{t-2}\\ \bar{{\bm{u}}}_{t-1}\otimes\bar{{\bm{u}}}_{t-2}\otimes{\bm{u}}_{t-3}\\ \vdots\\ \bar{{\bm{u}}}_{t-1}\otimes\bar{{\bm{u}}}_{t-2}\otimes\cdots\otimes{\bm{u}}_{t-L}\end{bmatrix},\qquad\tilde{{\bm{w}}}_{t}=\begin{bmatrix}{\bm{z}}_{t}\\ {\bm{w}}_{t-1}\\ {\bm{w}}_{t-2}\\ {\bm{w}}_{t-3}\\ \vdots\\ {\bm{w}}_{t-L}\end{bmatrix}$, (3.2)
$\text{and}\quad{\bm{F}}$ $=\begin{bmatrix}{\bm{I}}_{m}&{\bm{C}}&{\bm{C}}({\bm{u}}_{t-1}\circ{\bm{A}})&{\bm{C}}\prod_{\ell=1}^{2}({\bm{u}}_{t-\ell}\circ{\bm{A}})&\ldots&{\bm{C}}\prod_{\ell=1}^{L-1}({\bm{u}}_{t-\ell}\circ{\bm{A}})\end{bmatrix}$.
@@
${\bm{G}}$ $=\left[\begin{array}[]{c c c c c}{\bm{D}}&{\bm{G}}_{1}&{\bm{G}}_{2}&\cdots&{\bm{G}}_{L}\end{array}\right]\in\mathbb{R}^{m\times d_{\tilde{u}}},\quad\text{with}\quad d_{\tilde{u}}=(p+1)^{L}+p-1,$ (3.3)
@@
With the dynamics written in the form of (3.4), it is natural to use the least squares estimation method to learn the Markov-like parameters from the observations $\{{\bm{y}}_{t},{\bm{u}}_{t}\}_{t=1}^{T}$ . More specifically, the (minimum norm) Least Squares Estimator (LSE) $\hat{{\bm{G}}}$ of $\bm{G}$} admits a closed form and can be defined as:
@@
Let $\delta\in(0,1)$, $T\geq 0$. Suppose Assumption 1 holds, and the inputs are sampled uniformly at random from a sphere of radius $\sqrt{p}$, that is, $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{p-1})$. Then the event:
@@ 5 Numerical Experiments
For our experiments, we consider a partially observed bilinear dynamical system (2.1) with n=5n{=}5 hidden states, input dimension p=2p{=}2, and output dimension m=2m{=}2. The dynamics matrices ${\bm{A}}_{0},{\bm{A}}_{1},{\bm{A}}_{2}$ are constructed with i.i.d. $\mathcal{N}(0,1)$ entries, and are scaled to have spectral radius $\rho({\bm{A}}_{0}){=}\rho_{0}$ and $\rho({\bm{A}}_{1}){=}\rho({\bm{A}}_{2}){=}\rho_{k}$, where $\rho_{0},\rho_{k}$ are hyper-parameters in our experiments. Similarly, ${\bm{B}},{\bm{C}}$ and $\bm{D}$} are generated with i.i.d. $\mathcal{N}(0,1/n)$ and $\mathcal{N}(0,1/m)$ entries, respectively. The noise processes $\{{\bm{w}}_{t}\}_{t=0}^{T}$, and $\{{\bm{z}}_{t}\}_{t=0}^{T}$ are chosen according to i.i.d. Gaussian distribution with zero mean and variances $\sigma^{2}=0.0001$. Lastly, the control inputs are either sampled uniformly at random from the sphere $\sqrt{p}\cdot\mathcal{S}^{p-1}$ or sampled i.i.d. from a Gaussian distribution $\mathcal{N}(0,{\bm{I}}_{p})$.
In Figure 1, we plot the estimation error $\|{{\bm{G}}-\hat{{\bm{G}}}}\|_{\textup{op}}^{2}$ over different values of $\rho_{0},\rho_{k},L$ and $T$. Each experiment is repeated $10$ times and we plot the mean and one standard deviation. Figure 1(a) and Figure 1(c) correspond to estimation with Gaussian inputs $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{p})$, whereas, Figure 1(b) and Figure 1(d) correspond to estimation with uniformly distributed inputs $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{d-1})$.
Figure 1 shows that the estimation error increases with $L$, because the number of Markov-like parameters increases exponentially in $L$. More interestingly, Figure 1 suggests that choosing the inputs $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{d-1})$ leads to more accurate estimation than $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{p})$ at any given $T$. This seems to highlight the role of hypercontractivity parameter $\gamma$ in our estimation problem. Recall that $\gamma=3$ for Gaussian inputs, whereas, $\gamma=\frac{3}{1+2/p}$ for the uniformly distributed input.
We observe double descent curves [38] for L=7L{=}7. This is because our regression problem is unregularized and has $(p+1)^{L}+p-1=2188$ unknown parameters, and the number of input-features ($\tilde{{\bm{u}}}_{t}$) is T7T-7. Hence, for $L=7$, we see the peak at $T=2195$, and the error decays smoothly after this point. Note that the peak occurs at $T-L=(p+1)^{L}+p-1$ (where the number of unknown parameters become equal to the number of input-features). For $L=\{5,6\}$, we do not see the double descent because we start at $T=1000$ which is greater than $(p+1)^{L}+p-1$.
(a) $\{{\bm{u}}_{t}\}_{t\geq 0}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,{\bm{I}}_{p})$, $\rho_{0}{=}0.4,\rho_{k}{=}0.2$
(b) $\{{\bm{u}}_{t}\}_{t\geq 0}{\overset{\text{i.i.d.}}{\sim}}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{d-1})$, $\rho_{0}{=}0.4,\rho_{k}{=}0.2$
(c) $\{{\bm{u}}_{t}\}_{t\geq 0}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,{\bm{I}}_{p})$, $\rho_{0}{=}0.6,\rho_{k}{=}0.4$
(d) $\{{\bm{u}}_{t}\}_{t\geq 0}{\overset{\text{i.i.d.}}{\sim}}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{d-1})$, $\rho_{0}{=}0.6,\rho_{k}{=}0.4$
Figure 1: We plot the estimation error $\|{{\bm{G}}-\hat{{\bm{G}}}}\|_{\textup{op}}^{2}$ over different values of $T$, $L$, $\rho({\bm{A}}_{0})$, $\rho({\bm{A}}_{1})$, $\rho({\bm{A}}_{2})$ while fixing n=5n{=}5, p=2p{=}2 and m=2m{=}2. Figure 1(a) and Figure 1(c) correspond to estimation with Gaussian inputs $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{p})$, whereas, Figure 1(b) and Figure 1(d) correspond to estimation with uniformly distributed inputs $\{{\bm{u}}_{t}\}_{t\geq 0}\overset{\text{i.i.d.}}{\sim}\mathrm{Unif}(\sqrt{p}\cdot\mathcal{S}^{d-1})$. Our plots show that the later input choice gives better estimation of $\bm{G}$} as compared to the first one. Moreover, the estimation error increases as $L$ and $\rho$ increases, whereas, it decreases as $T$ increases.
@@ 6 Conclusion and Future Direction
Acknowledgements
S.D. was partly supported by NSF CCF 2312774, NSF OAC-2311521, NSF IIS-2442137, a PCCW Affinito-Stewart Award, a Gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, and an AI2050 Early Career Fellowship program at Schmidt Sciences. M.F. was supported in part by awards NSF TRIPODS II 2023166, NSF CCF 2007036, NSF CCF 2212261, NSF CCF 2312775, and by an Amazon-UW Hub gift award. Y.J. was partially supported by the Knut and Alice Wallenberg Foundation Postdoctoral Scholarship Program at MIT-KAW 2022.0366.
@@ B.2 Hypercontractivity to Bounded Moments (Covariates)
Isotropic Covariates: To begin, note that $\operatorname{\mathbb{E}}[\tilde{{\bm{u}}}_{t}]=0$ due to Assumption 2. Next, we will show that ${\bm{\Sigma}}[\tilde{{\bm{u}}}_{t}]=\operatorname{\mathbb{E}}[\tilde{{\bm{u}}}_{t}\tilde{{\bm{u}}}_{t}^{\top}]={\bm{I}}_{d_{\tilde{u}}}$ as follows: Let $\tilde{{\bm{u}}}_{t}]_{0}={\bm{u}}_{t}$, and let $\tilde{{\bm{u}}}_{t}]_{i}$ be the $i$-th partition of $\tilde{{\bm{u}}}_{t}$ with $p(p+1)^{i-1}$ entries for $i=1,\dots,L$. Then, we have that $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{0}[\tilde{{\bm{u}}}_{t}]_{0}^{\top}\right]=\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{1}[\tilde{{\bm{u}}}_{t}]_{1}^{\top}\right]={\bm{I}}_{p}$. Similarly, for $i=2,\dots,L$, we have
@@
for all $2\leq i<j$. Similarly, we also have $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{1}[\tilde{{\bm{u}}}_{t}]_{j}^{\top}\right]=\operatorname{\mathbb{E}}\left[{\bm{u}}_{t-1}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes{\bm{u}}_{t-j})^{\top}\right]=0$, for all $j=2,\dots,L$. Lastly, $\tilde{{\bm{u}}}_{t}]_{0}$ is independent of $\tilde{{\bm{u}}}_{t}]_{j}$, for all $j=1,\dots,L$, due to Assumption 2, hence, $\operatorname{\mathbb{E}}\left[[\tilde{{\bm{u}}}_{t}]_{0}[\tilde{{\bm{u}}}_{t}]_{j}^{\top}\right]=0$. Combining (B.2) and (B.3), we get
@@
$\bullet$ Quadratic Terms II: For $i\neq j$, consider the expectation $\operatorname{\mathbb{E}}\left[({\bm{v}}_{i}^{\top}[\tilde{{\bm{u}}}_{t}]_{i})^{2}({\bm{v}}_{j}^{\top}[\tilde{{\bm{u}}}_{t}]_{j})^{2}\right]$. In order to upper bound this term, we first upper bound an intermediate term as follows: Let $\tilde{\tilde{{\bm{u}}}}_{t}]_{k}$ denote the $k$-th partition of $\tilde{\tilde{{\bm{u}}}}_{t}$ defined in (B.1), for $k=1,\dots,L$. Let ${\bm{q}}_{k},{\bm{q}}^{\prime}_{k}$ be two arbitrary vectors in ${\bm{R}}^{(p+1)^{k}}$. We will use the mathematical induction to prove the following intermediate result under Assumption 2,
@@
$\qquad\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\left(q_{11}+{\bm{q}}_{12}^{\top}{\bm{u}}_{t-1}\right)^{2}\left(q_{11}^{\prime}+{\bm{q}}_{12}^{\prime\top}{\bm{u}}_{t-1}\right)^{2}\right],$
@@
$\qquad\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\big[\operatorname{\mathbb{E}}\big[\left({\bm{u}}_{t-i}^{\top}{\bm{V}}_{i}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes\bar{{\bm{u}}}_{t-i+1})\right)^{2}\left((\bar{{\bm{u}}}_{t-i}^{\top}\otimes\cdots\otimes{\bm{u}}_{t-j}^{\top}){\bm{V}}_{j}(\bar{{\bm{u}}}_{t-1}\otimes\cdots\otimes\bar{{\bm{u}}}_{t-i+1})\right)^{2}$
@@
$\qquad\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{k=1}^{(p+1)^{i-1}}\operatorname{\mathbb{E}}\big[\|{{\bm{V}}_{i}^{\top}{\bm{u}}_{t-i}}\|_{\ell_{2}}^{2}\big({\bm{v}}_{j2k}^{\top}({\bm{u}}_{t-i}\otimes\bar{{\bm{u}}}_{t-i-1}\otimes\cdots\otimes{\bm{u}}_{t-j})\big)^{2}\big],$
$\qquad\stackrel{{\scriptstyle\text{(b)}}}{{=}}\sum_{k=1}^{(p+1)^{i-1}}\operatorname{\mathbb{E}}\big[\|{{\bm{V}}_{i}^{\top}{\bm{u}}_{t-i}}\|_{\ell_{2}}^{2}\big((\bar{{\bm{u}}}_{t-i-1}\otimes\cdots\otimes{\bm{u}}_{t-j})^{\top}{\bm{V}}_{j2k}{\bm{u}}_{t-i}\big)^{2}\big],$
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[({\bm{v}}_{1}^{\top}{\bm{u}}_{t-1})^{2}\|{{\bm{v}}_{j1}+{\bm{V}}_{j2}{\bm{u}}_{t-1}}\|_{\ell_{2}}^{2}\right],$
@@
$\bullet$ Quartic Terms: Consider the expectation $\operatorname{\mathbb{E}}\left[({\bm{v}}_{i}^{\top}[\tilde{{\bm{u}}}_{t}]_{i})^{4}\right]$. In order to upper bound this term, we first upper bound an intermediate term as follows: Recall the definition of $\tilde{\tilde{{\bm{u}}}}_{t}$ from (B.1), and let $\tilde{\tilde{{\bm{u}}}}_{t}]_{k}$ denote its $k$-th partition with $(p+1)^{k}$ entries, for $k=1,\dots,L$. For any ${\bm{q}}_{k}\in{\bm{R}}^{(p+1)^{k}}$, we will use the mathematical induction to prove that
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\left(\bar{{\bm{u}}}_{t-k}^{\top}{\bm{Q}}_{k}[\tilde{\tilde{{\bm{u}}}}_{t}]_{k-1}\right)^{4}{~\big|~}{\bm{u}}_{t-k}\right]\right],$
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\left({\bm{u}}_{t-k}^{\top}{\bm{V}}_{k}[\tilde{\tilde{{\bm{u}}}}_{t}]_{k-1}\right)^{4}{~\big|~}{\bm{u}}_{t-k}\right]\right],$